# Larger Is Not Always Better: Exploring Small Open-source Language Models in Logging Statement Generation

### v2

@@ Abstract.
Developers use logging statements to create logs that document system behavior and aid in software maintenance. As such, high-quality logging is essential for effective maintenance; however, manual logging often leads to errors and inconsistency. Recent methods emphasize using large language models (LLMs) for automated logging statement generation, but these present privacy and resource issues, hindering their suitability for enterprise use. This paper presents the first large-scale empirical study evaluating small open-source language models (SOLMs) for automated logging statement generation. We evaluate four prominent SOLMs using various prompt strategies and parameter-efficient fine-tuning techniques, such as Low-Rank Adaptation (LoRA) and Retrieval-Augmented Generation (RAG). Our results show that fine-tuned SOLMs with LoRA and RAG prompts, particularly Qwen2.5-coder-14B, outperform existing tools and LLM baselines in predicting logging locations and generating high-quality statements, with robust generalization across diverse repositories. These findings highlight SOLMs as a privacy-preserving, efficient alternative for automated logging.
@@ 1. Introduction
Logs are textual records generated during software execution to capture runtime events, states, and contextual information (Yu et al., 2023b). A typical log statement consists of three components: a verbosity level, logging variables, and logging texts (He et al., 2021; Jiang et al., 2024b). In particular, as the example shown below, the logging level (e.g., debug) reflects the event’s severity; the logging variables (e.g., terminalState) hold critical run-time data about system states; meanwhile, the logging text (e.g., Stopping the checkpoint services with state) offers a static explanation of the system’s actions.
@@
High-quality logs provide actionable insights for developers to diagnose failures, optimize system behavior, and ensure reliability. However, the absence or inadequacy of logging statements can severely hinder downstream tasks such as anomaly detection (Zhang et al., 2022; Li et al., 2020b; Liu et al., 2023) and failure diagnosis (Yu et al., 2023a; Jiang et al., 2025; Huang et al., 2024b), leading to prolonged debugging cycles and increased maintenance costs. Consequently, the strategic placement and content of log statements directly influence the effectiveness of software maintenance and evolution (Soldani and Brogi, 2022; Huang et al., 2024a).
@@
To address these challenges, at the early stage, researchers have explored automated logging techniques by decomposing the problem into sub-tasks, such as where-to-log (identifying code locations for logging) (Zhao et al., 2017; Li et al., 2020a), what-to-log (generating static content and dynamic variables) (Ding et al., 2022; Liu et al., 2019), and log-level suggestion (Heng et al., 2024; Li et al., 2021a; Liu et al., 2022). However, these fragmented approaches lack integration into a unified, end-to-end framework for generating complete logging statements. Recent advances in pre-trained language models (LMs) have opened new avenues for automated logging. LANCE (Mastropaolo et al., 2022) and LEONID (Mastropaolo et al., 2024) pioneered the use of sequence-to-sequence models (e.g., T5 (Raffel et al., 2020)) to generate logging statements directly from code contexts. Subsequent tools, such as Fastlog (Xie et al., 2024), Unilog (Xu et al., 2024), and SCLogger (Li et al., 2024b), further leveraged the large language model (LLM) to improve logging quality. In particular, Unilog and SCLogger adopted prompt-based methods with LLMs such as GPT-3.5 and Codex, achieving state-of-the-art performance by harnessing the code comprehension and natural language generation capabilities of LLMs.
@@
RQ4: Can SOLMs generalize logging statement generation across diverse code repositories? Logging practices vary across projects. Evaluating SOLM generalization ability on diverse repositories ensures their applicability in real-world development environment.
Result. SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories. Similar logging practices, such as those shared among Apache open-source projects, significantly improve the cross-project generalization ability of SOLMs.
@@
Paper Orgenizations. Section 2 discusses the background. Section 3 describes the experimental design of our study. Section 4 presents the experimental results. Section 5 introduces the advantages of using SOLMs for automated logging and potential future work directions. Section 6 discusses threats to validity. Section 7 introduces the related work. Section 8 concludes the paper.
@@ 2.2. Large Language Models
The evolution of language models in recent times can be divided into three transformative phases. Initially, there were neural language models (NLMs), followed by the phase of pre-trained language models (PLMs), and the current era sees the prominence of LLMs. Pre-trained models such as CodeT5 (Wang et al., [n. d.]) and PLBART (Ahmad et al., 2021) have achieved noteworthy success in software engineering applications primarily due to task-specific pre-training processes. However, large language models have brought about a revolutionary change in the field due to their immense parameter counts, often exceeding 10 billion, and their comprehensive pre-training data. These models, unlike their pre-training predecessors, exhibit emergent capabilities that allow them to achieve robust performance across a wide range of tasks without necessitating pre-training tailored to specific tasks (Geng et al., 2024). This quality substantially diminishes the requirement for resource-heavy training sessions. Within the realm of software engineering, large language models are mainly categorized into two groups. Unified large language models, such as GPT-4o and LLaMA, which are designed to integrate natural language and code corpora, whereas code-specific large language models, like StarCoder (Li et al., 2023a) and CodeLlama (Roziere et al., 2023), are developed for specialization in tasks centered around coding.
Current methodologies for leveraging LLMs fall into two paradigms: prompt-based and tuning-based. Prompt-based methods exploit the zero-shot or few-shot capabilities of massive LLMs (e.g., GPT-4) through carefully engineered prompts, avoiding the need for explicit training data. They use techniques such as in-context learning (ICL) and chain-of-thought (COT). Conversely, fine-tuning approaches focus on parameter-efficient adaptation (e.g., prompt tuning, prefix tuning, and low-rank adaptation) to tailor smaller-scale LLMs for domain-specific tasks. These techniques freeze base model parameters while training minimal additional components, achieving performance comparable to full-parameter tuning at reduced computational costs.
@@ 2.3. LLM Applications in Software Engineering Task
In certain tasks, approaches that utilize prompts have been shown to yield superior outcomes (Imani et al., 2024; Jiang et al., 2024a). Pan et al., for instance, studied the efficiency of LLMs in the context of code translation (Pan et al., 2024). Within the range of models assessed, which encompassed both SOLMs and GPT-4, the top-performing SOLM, known as StarCoder, managed to reach a successful translation rate of 14.5%, in contrast to the success rate of 47.3% achieved by GPT-4. In contrast, SOLMs have been shown to achieve comparable outcomes in specific domains. For instance, Tian et al. (Tian et al., 2024) observed an F1-score of 86.58% using UniXCoder for the task of detecting equivalent mutants, providing a notable contrast to the performance of GPT-4, which recorded an F1-score of approximately 55.90%. Furthermore, when employing Vicuna 13B in conjunction with the innovative LogPrompt strategy, the performance was found to be on par with that of GPT-4, as reported in (Liu et al., 2024b).
@@ 3.1. Dataset Preparation
To evaluate the performance of automated logging, we select the most recently proposed AL-Bench (Tan et al., [n. d.]). This dataset, containing 42,224 logging statements from 10 widely used projects, was specifically constructed using stringent criteria (e.g., ¿=10k stars, ¿=1k logging statements, ¿=500 log-related issues per project) to ensure the inclusion of well-maintained repositories where developers prioritize high-quality logging statements. Furthermore, AL-Bench intentionally incorporates projects from diverse domains, such as database management, task scheduling, distributed systems, messaging, and IoT platforms, to ensure comprehensive coverage of various real-world logging requirements and practices, making it a robust benchmark. Measures like using recent project versions and standardized code formatting were also employed during its creation to mitigate potential data contamination risks from pre-training corpora. Table 1 provides further statistical details on these source repositories.
@@
To align with the research objectives of this study and accommodate the requirements of certain baseline models, we performed several adjustments to the original dataset. Firstly, recognizing that some baseline models have an input token limit, we filtered out data instances where the input code exceeded 512 tokens. Secondly, the original construction of AL-Bench could generate multiple data points from a single Java function if it contained multiple logging statements, with each data point representing one specific automated logging case. To prevent potential data leakage, where highly similar code snippets from the same source file might inadvertently appear across different dataset splits (e.g., fine-tuning and testing), we implemented a file-level splitting strategy. This approach ensures that all data instances originating from the same Java file are strictly allocated to only one of the fine-tuning, validation, or test sets. After applying these pre-processing steps, we obtained a final dataset comprising 33,224 instances. We then partitioned this dataset into fine-tuning, validation, and test sets, targeting an 8:1:1 ratio. Because our file-level splitting strategy required keeping all instances from a single file within the same set, the resulting distribution was approximate. The final fine-tuning set contains 26,713 instances, the validation set contains 3,508 instances, and the test set contains 3,003 instances.
@@ 3.3. Baselines
To evaluate SOLMs performance, we select the baselines by conducting a literature review of relevant papers published in SE venues. From this, we find the following methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2024), UniLog (Xu et al., 2024), and FastLog (Xie et al., 2024). Additionally, we examine several LLM baselines, including general-purpose LLMs (Claude3.7-sonnet, GPT4o, LLaMA3.1-405b), and a code-specific LLM (Deepseek-coder-v3). In relation to the methodologies for prompting, we derive our approach from the study (Gao et al., 2024) and incorporate four different strategies: instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG) and chain of thought (CoT). The instruction prompting strategy involves directly prompting LLMs to generate logging statements using identical inputs as those provided to SOLMs, without any supplementary data. The ICL approach consists of providing one random example before the main query to assist the model in grasping the nature of the task more effectively. In the RAG approach, rather than selecting example randomly, we retrieve precise examples from the valid dataset to accommodate various input samples. We specifically employ the BM25 algorithm to identify the case within the valid dataset that most closely resembles the query sample in terms of similarity. The details of the prompts are shown in Figure 3.
@@ 3.4. Strategies for Parameter-Efficient Fine-Tuning
Lora (Hu et al., 2022) provides a distinct PEFT mechanism based on the hypothesis that the change in weights during model adaptation has a low intrinsic rank. Instead of adding prefix or prompt tokens, LoRA freezes the original pre-trained weights $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ of selected layers (commonly the query, key, value, and output projection matrices in self-attention, and sometimes feed-forward layers) and injects trainable, rank-decomposition matrices in parallel. Specifically, the weight update $\Delta\mathbf{W}$ is approximated by the product of two smaller, low-rank matrices: $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{r\times k}$, where the rank $r$ is a crucial hyperparameter significantly smaller than the original dimensions ($r\ll\min(d,k)$). The modified forward pass for an input $\mathbf{x}$ computes the output as the sum of the original path and the adapter path: $\mathbf{h}_{\text{adapted}}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\alpha(\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}})\mathbf{x}$, where $\alpha$ is often a constant scaling factor (like $\alpha/r$). During fine-tuning, only the parameters of $\mathbf{W}_{\text{down}}$ (typically initialized randomly) and $\mathbf{W}_{\text{up}}$ (typically initialized to zero) are optimized. A significant advantage of LoRA is that the learned adapter weights can be mathematically merged with the original weights ($\mathbf{W}=\mathbf{W}_{0}+\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}}$) after training, resulting in a single weight matrix per adapted layer and incurring zero additional inference latency compared to the original model, while still offering substantial parameter savings during training and allowing easy task switching by loading different adapter pairs.

### v3

@@ Abstract.
Developers use logging statements to create logs that document system behavior and aid in software maintenance. As such, high-quality logging is essential for effective maintenance; however, manual logging often leads to errors and inconsistency. Recent methods emphasize using large language models (LLMs) for automated logging statement generation, but these present privacy and resource issues, hindering their suitability for enterprise use. This paper presents the first large-scale empirical study evaluating small open-source language models (SOLMs) for automated logging statement generation. We evaluate four prominent SOLMs using various prompt strategies and parameter-efficient fine-tuning techniques, such as Low-Rank Adaptation (LoRA) and Retrieval-Augmented Generation (RAG). Our results show that fine-tuned SOLMs with LoRA and RAG prompts, particularly Qwen2.5-coder-14B, outperform existing tools and LLM baselines (e.g., Claude3.7 sonnet and GPT4o) in predicting logging locations and generating high-quality statements, with robust generalization across diverse repositories. These findings highlight SOLMs as a privacy-preserving, efficient alternative for automated logging.
@@ 1. Introduction
Logs are textual records generated during software execution to capture runtime events, states, and contextual information (Zhong, 2025). A typical log statement consists of three components: a verbosity level, logging variables, and logging texts (He et al., 2021). In particular, as the example shown below, the logging level (e.g., debug) reflects the event’s severity; the logging variables (e.g., terminalState) hold critical run-time data about system states; meanwhile, the logging text (e.g., Stopping the checkpoint services with state) offers a static explanation of the system’s actions.
@@
High-quality logs provide actionable insights for developers to diagnose failures, optimize system behavior, and ensure reliability. However, the absence or inadequacy of logging statements can severely hinder downstream tasks such as anomaly detection (Zhang et al., 2022) and failure diagnosis (Jiang et al., 2025; Huang et al., 2024), leading to prolonged debugging cycles and increased maintenance costs. Consequently, the strategic placement and content of log statements directly influence the effectiveness of software maintenance and evolution (Soldani and Brogi, 2022; Gu et al., 2025).
@@
To address these challenges, at the early stage, researchers have explored automated logging techniques by decomposing the problem into sub-tasks, such as where-to-log (identifying code locations for logging) (Zhao et al., 2017; Li et al., 2020), what-to-log (generating static content and dynamic variables) (Ding et al., 2022; Liu et al., 2019), and log-level suggestion (Heng et al., 2024; Li et al., 2021a; Liu et al., 2022). However, these fragmented approaches lack integration into a unified, end-to-end framework for generating complete logging statements. Recent advances in pre-trained language models (LMs) have opened new avenues for automated logging. LANCE (Mastropaolo et al., 2022) and LEONID (Mastropaolo et al., 2024) pioneered the use of sequence-to-sequence models (e.g., T5 (Raffel et al., 2020)) to generate logging statements directly from code contexts. Subsequent tools, such as Fastlog (Xie et al., 2024), Unilog (Xu et al., 2024), and SCLogger (Li et al., 2024b), further leveraged the large language model (LLM) to improve logging quality. In particular, Unilog and SCLogger adopted prompt-based methods with LLMs such as GPT-3.5 and Codex, achieving state-of-the-art performance by harnessing the code comprehension and natural language generation capabilities of LLMs.
@@
RQ4: How robust are SOLMs in handling diverse real-world scenarios for logging statement generation? We assess robustness by evaluating their generalization across diverse repositories and analyzing the performance impact of varying code lengths.
Result. SOLMs demonstrate strong robustness, generalizing effectively to unseen repositories. Performance varies with code length, revealing a trade-off between positioning accuracy and content quality. We also find that similar logging practices across projects significantly boost generalization.
Table 1. Summarization of Key Findings and Implications in This Paper.
Key Findings Key Implications & Actionable Advice
>>>> RQ1 & RQ2: Prompting and Tuning Strategies
[…]
>>>> RQ3: Baseline Comparison
[…]
>>>> RQ4: Robustness and Generalization
[…]
In Table 1, we summarize the thirteen findings and three implications with actionable advice.
@@
Paper Organizations. Section 2 discusses the background. Section 3 describes the experimental design of our study. Section 4 presents the experimental results. Section 5 introduces a case study, some implications, the advantages of using SOLMs for automated logging, and potential future work directions. Section 6 discusses threats to validity. Section 7 introduces the related work. Section 8 concludes the paper.
@@ 2.2. Large Language Models
The evolution of language models in recent times can be divided into three transformative phases. Initially, there were neural language models (NLMs), followed by the phase of pre-trained language models (PLMs), and the current era sees the prominence of LLMs. Pre-trained models such as CodeT5 (Wang et al., [n. d.]) and PLBART (Ahmad et al., 2021) have achieved noteworthy success in software engineering applications primarily due to task-specific pre-training processes. However, LLMs have brought about a revolutionary change in the field due to their immense parameter counts, often exceeding 10 billion, and their comprehensive pre-training data. These models, unlike their pre-training predecessors, exhibit emergent capabilities that allow them to achieve robust performance across a wide range of tasks without necessitating fine-tuning tailored to specific tasks (Geng et al., 2024). This quality substantially diminishes the requirement for resource-heavy training sessions. Within the realm of software engineering, LLMs are mainly categorized into two groups. Unified large language models, such as GPT-4o and LLaMA, which are designed to integrate natural language and code corpora, whereas code-specific large language models, like StarCoder (Li et al., 2023a) and CodeLlama (Roziere et al., 2023), are developed for specialization in tasks centered around coding.
Methodologies for leveraging these models typically follow two distinct paradigms based on model scale and accessibility. The first is the prompt-based paradigm, which exploits the zero-shot or few-shot capabilities of massive-scale LLMs (e.g., GPT-4). This approach relies on carefully engineered prompts, using techniques like in-context learning (ICL) and chain-of-thought (CoT) to guide the model without updating its parameters. This paradigm is characteristic of the largest models, which often pose significant resource challenges and are typically accessed via APIs.
A contrasting paradigm involves the tuning-based adaptation of smaller-scale, open-source models through PEFT techniques like Low-Rank Adaptation (LoRA). This approach has given rise to a focus on what this paper terms SOLMs, which are designed to balance performance with practical deployability. While the exact parameter threshold for such models is an evolving topic, they are often distinguished from massive-scale LLMs by a specific size cap, such as under 10 billion parameters (Yang et al., 2024b; Fu et al., 2023). In this study, we adopt a hardware-centric threshold and define SOLMs as models with fewer than 14 billion parameters. This practical boundary reflects the limit of what can be efficiently fine-tuned and deployed on a single, widely available high-end GPU (e.g., an NVIDIA A100 80G), a rationale similarly employed for defining lightweight models in other software engineering contexts (Yang et al., 2024b). This distinction is critical, as SOLMs represent a class of models that can be leveraged by enterprises to create specialized, cost-effective, and privacy-preserving solutions for tasks like automated logging, circumventing the resource demands of their larger counterparts.
@@ 2.3. LLM Applications in Software Engineering Task
In certain tasks, approaches that utilize prompts have been shown to yield superior outcomes (Imani et al., 2024; Jiang et al., 2024). Pan et al., for instance, studied the efficiency of LLMs in the context of code translation (Pan et al., 2024). Within the range of models assessed, which encompassed both SOLMs and GPT-4, the top-performing SOLM, known as StarCoder, managed to reach a successful translation rate of 14.5%, in contrast to the success rate of 47.3% achieved by GPT-4. In contrast, SOLMs have been shown to achieve comparable outcomes in specific domains. For instance, Tian et al. (Tian et al., 2024) observed an F1-score of 86.58% using UniXCoder for the task of detecting equivalent mutants, providing a notable contrast to the performance of GPT-4, which recorded an F1-score of approximately 55.90%. Furthermore, when employing Vicuna 13B in conjunction with the innovative LogPrompt strategy, the performance was found to be on par with that of GPT-4, as reported in (Liu et al., 2024b).
@@ 3.1. Dataset Preparation
To evaluate the performance of automated logging, we selected AL-Bench (Tan et al., [n. d.]), the most recently proposed large-scale benchmark for this task. Our choice was deliberate, as AL-Bench offers several critical advantages over prior datasets that are essential for a rigorous evaluation of SOLMs and LLMs.
First, AL-Bench is specifically designed to evaluate the end-to-end task of logging statement generation, encompassing both the placement (where-to-log) and content creation (what-to-log) aspects. Second, the benchmark was constructed using stringent quality criteria (e.g., ¿=10k stars, ¿=1k logging statements, ¿=500 log-related issues per project), sourcing data exclusively from 10 popular, well-maintained Java projects across a diverse set of domains, including database management, task scheduling, and IoT platforms. This ensures that the ground truth logging statements are of high quality and that our evaluation is representative of real-world practices.
Most importantly, AL-Bench was developed with a keen awareness of the challenges posed by LLMs, particularly the threat of data leakage. The benchmark’s creators employed specific mitigation strategies to minimize the likelihood that its content was included in the pre-training corpora. For instance, all code snippets were processed with the Google-Java-Format tool and wrapped in a generic class, a method designed to alter the code’s original structure and appearance. This thoughtful step, combined with the use of the most recent project versions, further reduces the risk that a model’s performance might be inflated by prior exposure to the test data.
The thoughtful design of AL-Bench provides a robust foundation for evaluating modern language models. To ensure a consistent and fair comparison across all methods, we adopted AL-Bench as the sole benchmark for our experiments. This unified approach allows us to rigorously evaluate our fine-tuned models alongside existing state-of-the-art tools within the same controlled environment. Table 2 provides further statistical details on these source repositories.
@@
To construct the primary dataset used for our main experiments (RQ1-RQ3, RQ4.1), we performed several pre-processing steps on the original AL-Bench dataset. Firstly, to ensure a fair comparison and accommodate the input constraints of certain baseline models, we established a maximum input length of 512 tokens. Data instances with code snippets exceeding this threshold were set aside for a separate analysis. Secondly, the original construction of AL-Bench could generate multiple data points from a single Java function if it contained multiple logging statements, with each data point representing one specific automated logging case. To prevent potential data leakage, where highly similar code snippets from the same source file might inadvertently appear across different dataset splits (e.g., fine-tuning and testing), we implemented a file-level splitting strategy. This approach ensures that all data instances originating from the same Java file are strictly allocated to only one of the fine-tuning, validation, or test sets. After applying these pre-processing steps, we obtained a final dataset comprising 33,224 instances. We then partitioned this dataset into fine-tuning, validation, and test sets, targeting an 8:1:1 ratio. Because our file-level splitting strategy required keeping all instances from a single file within the same set, the resulting distribution was approximate. The final fine-tuning set contains 26,713 instances, the validation set contains 3,508 instances, and the test set contains 3,003 instances.
In addition, to specifically investigate the impact of code length in RQ4.2, we utilized the previously set-aside longer code snippets to construct a specialized test set. This test set is also going through the above pre-processing steps. The detailed motivation and use of this set are further elaborated in the approach part of RQ4.2.
@@ 3.3. Baselines
To evaluate SOLMs performance, we select the baselines by conducting a literature review of relevant papers published in SE venues. From this, we find the following methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2024), UniLog (Xu et al., 2024), FastLog (Xie et al., 2024), and SCLogger (Li et al., 2024b). Additionally, we examine several LLM baselines, including general-purpose LLMs (Claude3.7-sonnet, GPT4o, LLaMA3.1-405b), and a code-specific LLM (Deepseek-coder-v3).
In relation to the methodologies for prompting, we derive our approach from the study (Gao et al., 2024) and incorporate four different strategies: instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG) and chain of thought (CoT). The instruction prompting strategy involves directly prompting LLMs to generate logging statements using identical inputs as those provided to SOLMs, without any supplementary data. The ICL approach consists of providing one random example before the main query to assist the model in grasping the nature of the task more effectively. In the RAG approach, we enhance the prompt by providing the most similar example to guide the model. Instead of selecting an example randomly as in ICL, we retrieve the most similar instance from our validation set to serve as a one-shot demonstration. This process leverages the validation set as a dedicated knowledge pool. Specifically, for each input sample from the test set, we employ the BM25 algorithm (Robertson et al., 2009) to identify and retrieve the single most similar code snippet from the validation set. This retrieved example is then prepended to the main query within the prompt. Finally, for the CoT strategy, we employ a zero-shot approach that explicitly instructs the model to ”reason step-by-step” about the purpose, placement, content, and level of the logging statement before providing the final code output. The details of the prompts are shown in Figure 3.
@@ 3.4. Strategies for Parameter-Efficient Fine-Tuning
LoRA (Hu et al., 2022) provides a distinct PEFT mechanism based on the hypothesis that the change in weights during model adaptation has a low intrinsic rank. Instead of adding prefix or prompt tokens, LoRA freezes the original pre-trained weights $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ of selected layers (commonly the query, key, value, and output projection matrices in self-attention, and sometimes feed-forward layers) and injects trainable, rank-decomposition matrices in parallel. Specifically, the weight update $\Delta\mathbf{W}$ is approximated by the product of two smaller, low-rank matrices: $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{r\times k}$, where the rank $r$ is a crucial hyperparameter significantly smaller than the original dimensions ($r\ll\min(d,k)$). The modified forward pass for an input $x$ computes the output by merging the original and adapter paths: $h_{\text{adapted}}=\mathbf{W}_{0}x+\Delta\mathbf{W}x=\mathbf{W}_{0}x+\alpha(\mathbf{W}_{\text{down}}\mathbf{W}_{\text{up}})x$. In this formulation, $\alpha$ is a scaling hyperparameter. In some implementations, this scaling factor is set in proportion to the rank $r$ (e.g., as $\frac{\alpha}{r}$), which helps stabilize the adaptation process by ensuring the magnitude of the update remains consistent when experimenting with different values of $r$. During fine-tuning, only the parameters of $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{up}}$ are optimized. A significant advantage of LoRA is that the learned adapter weights can be mathematically merged with the original weights after training, resulting in a single weight matrix per adapted layer and incurring zero additional inference latency compared to the original model, while still offering substantial parameter savings during training and allowing easy task switching by loading different adapter pairs.